OpenStudy (anonymous):

What is the solution set for |–2x + 16| – 9 = –3 x = –5 and x = –11 x = 5 and x = 11 x = –5 and x = 11 x = 5 and x = –11

OpenStudy (anonymous):

|-2x + 16| = 6 -2x + 16 = 6 or -2x + 16 = -6

OpenStudy (anonymous):

thanks